I need to know an alternate way of expressing a logarithm
log m = log a + 3 log l

Answer by longjonsilver(2297)

log1 in any base is zero, so we straight away get
log m = log a
log m - log a = 0
log (m/a) = 0

Now, you do not say which base these logs are in. I will assume it is base 10, so we now raise everything to that base, giving:

(m/a) = 10^0
(m/a) = 1
therefore, m = a

This answer is "obvious" from the very first line --> log m = log a, where we have log of a number equals log of another number. Only way this is possible is if the 2 numbers are one in the same.

jon.